# LLM Collaboration with Multi-Agent Reinforcement Learning

### v2

@@ Abstract
A large amount of work has been done in Multi-Agent Systems (MAS) for modeling and solving problems with multiple interacting agents. However, most LLMs are pretrained independently and not specifically optimized for coordination. Existing LLM fine-tuning frameworks rely on individual rewards, which require complex reward designs for each agent to encourage collaboration. To address these challenges, we model LLM collaboration as a cooperative Multi-Agent Reinforcement Learning (MARL) problem. We develop a multi-agent, multi-turn algorithm, Multi-Agent Group Relative Policy Optimization (MAGRPO), to solve it, building on current RL approaches for LLMs as well as MARL techniques. Our experiments on coding collaboration demonstrate that fine-tuning MAS with MAGRPO enables agents to generate high-quality responses efficiently through effective cooperation. Our approach opens the door to using other MARL methods for LLMs and highlights the associated challenges.
1 Introduction
@@
Leveraging billions of parameters and extensive pre-training on large-scale datasets, state-of-the-art LLMs have demonstrated remarkable capabilities across diverse domains Grattafiori et al. (2024); Achiam et al. (2023); Anil et al. (2025). To adapt to specific applications or align with human preferences, fine-tuning has emerged as a critical secondary training stage. Compared to supervised fine-tuning, Reinforcement Learning (RL) enables more generalizable learning for complex, multi-turn tasks through human-aligned reward design, making it an important technique for fine-tuning Ouyang et al. (2022); Guo et al. (2025); Ziegler et al. (2020).
@@
Likewise, Multi-Agent Systems (MAS) have been extensively studied over the past decades, with substantial progress in modeling and solving problems involving multiple agents Littman (1994); Shoham and Leyton-Brown (2009); Stone and Veloso (2000). In particular, advances in cooperative MAS have demonstrated strong potential for enabling effective collaboration in distributed settings, such as games, robotics, and traffic control Samvelyan et al. (2019); Vinyals et al. (2017); Berner et al. (2019); Amato et al. (2016); Wiering (2000); Liu et al. (2022). These developments motivate the application of MAS principles and techniques to LLM collaboration, where multiple LLMs working together can solve more complex tasks in a more robust and efficient manner.
@@
There has been some recent work on coordinating multiple LLMs. Some approaches implement coordination at the inference stage, enabling agents to interact through debate, discussion, or verification Du et al. (2023); Wu et al. (2023a); Lifshitz et al. (2025). These methods operate at the prompt level, with fixed models that are not tuned toward coordination-centric objectives. The agents may have conflicting answers or spread incorrect information to other participants, limiting performance Cemri et al. (2025); Estornell and Liu (2024). Moreover, the design of effective prompts remains difficult and unclear. Other approaches fine-tune agents independently with individual or role-conditioned rewards. However, they require carefully curated rewards for each individual or role Slumbers et al. (2024); Liu et al. (2025a); Subramaniam et al. (2025), and, as independent learning methods, lack convergence guarantees Tan (1993).
@@
In this paper, we model LLM collaboration as a cooperative MARL problem Albrecht et al. (2024) and formalize it as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) Oliehoek and Amato (2016). In LLM collaboration, multiple trainable LLMs generate responses synchronously based on their individual prompts. The external environment evolves according to the joint responses until the dialog ends. This general model allows a wide range of problems to be modeled and solved using versions of MARL algorithms. Following the efficient practice of Group Relative Policy Optimization (GRPO) Guo et al. (2025), we propose Multi-Agent GRPO (MAGRPO) that trains LLMs in a multi-turn setting. MAGRPO leverages centralized group-relative advantages for joint optimization, while preserving decentralized execution for each agent. Our method builds off of state-of-the-art LLM approaches in GRPO and MARL approaches for centralized training and decentralized execution, such as MAPPO Yu et al. (2022). Our experiments show that MAGRPO develops various cooperation schemes, improving response efficiency with high quality.
@@
Our contributions can be summarized as follows: (i) We model the LLM collaboration as a cooperative MARL problem, where multiple LLMs cooperate to generate joint responses; (ii) We implement the MAGRPO algorithm, which optimizes agent cooperation through aligned rewards while maintaining decentralized execution to maintain efficiency; (iii) Our experiments demonstrate that fine-tuning with MAGRPO improves both response efficiency and quality in coding collaboration; (iv) We provide a detailed analysis of the limitations of existing approaches and outline open challenges in applying MARL to LLM collaboration.
2 Related Work
2.1 Test-Time Multi-Agent Interaction
@@
Recent work employs multiple agents with specialized roles interacting through diverse pipelines at test-time to enhance response quality. In multi-agent debate, agents iteratively formulate positions by reviewing other agents’ outputs, where the final decision or answer is determined by majority voting or a summarizer Du et al. (2023); Chan et al. (2023); Liang et al. (2024). Role-based approaches allocate tasks across specialized agents Wu et al. (2023a); Qian et al. (2024); Hong et al. (2024). For example, an agent may function as a verifier to assess the correctness of outputs Skreta et al. (2023); Lifshitz et al. (2025); Setlur et al. (2025), while another may act as a macro-planner to orchestrate workers’ responses. However, these multi-agent frameworks rely on prompt-level interactions among agents, often leading to ineffective communication and computational inefficiency. Moreover, the design of effective prompts and role assignment remains unclear, as prompts usually fail to reliably guide agent behavior, enforce role adherence, or support coherent coordination across tasks. These limitations motivate us to fine-tune LLMs in MAS to improve their cooperation.
2.2 Multi-Agent Fine-Tuning
@@
Recent work has explored fine-tuning LLMs to improve their performance across diverse domains, e.g., arithmetic reasoning, navigation, and hidden-role games Ma et al. (2025); Slumbers et al. (2024); Sarkar et al. (2025). These approaches typically employ individual rewards or rewards conditioned on specific roles Park et al. (2025); Liu et al. (2025a); Subramaniam et al. (2025). Such reward structures often require careful manual specification, and their underlying rationale is rarely well justified. The misaligned or conflicting incentives can hinder effective coordination. Moreover, these methods lack convergence guarantees, as each agent learns independently in a non-stationary environment where other agents are simultaneously updating their policies. In this paper, we focus on cooperative scenarios, where LLMs are jointly trained with interpretable, human-aligned rewards.
3 Cooperative MARL for LLM Collaboration
@@
Since LLMs can be viewed as a special class of agents, we leverage advances in MAS to facilitate their collaboration. We model LLM collaboration as a cooperative MARL problem and outline its unique challenges. We formalize this problem as a Dec-POMDP, as shown in Figure 1.
3.1 LLM Collaboration
LLM collaboration refers to the problems where LLMs cooperatively solve a class of tasks in MAS. The tasks are specified in natural language and provided to the agents as prompts. Each LLM agent generates a response synchronously based on its individual instructions. The set of these responses jointly forms a solution to the task.
@@ 3.1 LLM Collaboration
As discussed by a number of companies NVIDIA (2024); Anthropic (2024), a team of agents could be used to generate a complex codebase. The code would be difficult, costly, and time-consuming to generate with a single agent, but a group of LLMs could do so quickly and cheaply. None of these agents is self-interested, but they are trainable in a scheme such as the one discussed below. Using a joint reward allows agents to specialize as needed to complete the task without complex prompt or reward engineering.
3.2 Problem Formalization
We formalize collaboration among LLMs as a subclass of the cooperative MARL problem, considering LLM agents and the types of problems they are solving. This problem is a form of a Dec-POMDP Oliehoek and Amato (2016), which allows cooperation through a joint reward while preserving scalable decentralized control.
@@ 3.2 Problem Formalization
  • $\mathcal{S}$ denotes the full global state space. At turn $t$, a full state $s_{t}=(s_{t}^{\text{acc}},s_{t}^{\text{usr}})$ consists of parts that are accessible in the model and provided to the reward model $s_{t}^{\text{acc}}\in\mathcal{S}^{\text{acc}}$ (e.g., external models or systems), and the inaccessible user state $s_{t}^{\text{usr}}\in\mathcal{S}^{\text{usr}}$ which updates over time but isn’t maintainable. In Dec-POMDP, the state is not directly observable by the agents (LLMs).
@@
In Dec-POMDP, since the states are not directly observed, each agent maintains its local observation-action history $\mathbf{h}=\{h_{1},\cdots,h_{n}\}$ to infer information about state. A solution to a Dec-POMDP is a joint policy that maximizes the expected cumulative reward, $\boldsymbol{\pi}^{*}=\{\pi_{1}^{*},\cdots,\pi_{n}^{*}\}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{H-1}R(s^{\text{acc}}_{t},\mathbf{a}_{t})\right]$. A joint policy is a set of local policies $\pi_{i}$, which conditions on the local observation-action history $h_{i,t}=\{o_{i,0},a_{i,0},\ldots,o_{i,t}\}$.
RL methods for Dec-POMDPs have become a popular topic (e.g., Foerster et al. (2024); Lowe et al. (2020); Foerster et al. (2018); Rashid et al. (2018); Wang et al. (2021); Yu et al. (2022); Albrecht et al. (2024)) with methods successful at scaling to large state, action and observation spaces. Many methods use Centralized Training for Decentralized Execution (CTDE), where they use some centralized information during training (e.g., a centralized value function estimate) but are still able to execute in a decentralized manner when training is complete.
3.3 Challenges in LLM Collaboration
@@
3.3.1 Representations in Natural Language
Unlike traditional cooperative MARL agents, LLM agents operate over natural language, receiving instructions and generating responses as sequences of tokens. MARL approaches could model this problem at the token or prompt/response level. At the token level, the number of actions and observations is smaller, but the problem horizon can be very long. At the prompt/response level, the actions and observations space is much larger, but the horizon is much shorter. Moreover, token-level rewards are often uninformative, as both queries and responses must form coherent and semantically meaningful structures. As adopted in prior RL methods Ouyang et al. (2022); Guo et al. (2025); Rafailov et al. (2024), we model each LLM agent’s decision-making process as a direct mapping from input instructions to complete responses to enable efficient and stable training. Nevertheless, the best modeling and solution approaches remain an open question.
3.3.2 Training Paradigm
As mentioned above, many MARL methods use centralized training for decentralized execution (CTDE). Unfortunately, standard CTDE methods use centralized value models in the form of centralized critics Foerster et al. (2024); Lowe et al. (2020); Yu et al. (2022) or mixers in value decomposition methods Rashid et al. (2018); Wang et al. (2021). Such architectures allow additional information and coordination during training but do not scale well to very large action and observation spaces (such as those in our problem). Conversely, Decentralized Training and Execution (DTE) methods Amato (2025) train a set of models, one for each agent in a decentralized manner. DTE approaches are typically more scalable but do not use additional information during training (even when it is available). It is an open question which paradigm to use to maximize performance while maintaining scalability in the LLM collaboration problem. In this paper, we balance decentralized execution with centralized training using group-based Monte Carlo estimates. Experiments show the effectiveness of our approach on short-horizon tasks.
4 MAGRPO
@@
7:   Obtain a joint reward $r^{\mathcal{G}}_{t}$ from system
@@
12:   Estimate $\widehat{A}^{(g)}_{t}$, $\forall g\in\mathcal{G}$ for each branch $B$ according to Equation 1
13:   Calculate $J(\theta_{i})$, $\forall i\in\mathcal{I}$ for each branch $B$ according to Equation 2
@@
Algorithm 1 shows the procedure of MAGRPO. Given a dataset $\mathcal{D}$ containing task information (e.g., the descriptions of coding problems), $n$ LLMs are optimized, each with a policy parameterized by $\theta_{i}$ and guided by a reward model $R$. In each episode, a task is sampled from the given dataset $\mathcal{D}$, which is used to construct initial observations $\mathbf{o}_{0}=\{o_{1,0},\cdots,o_{n,0}\}$ and histories $\mathbf{h}_{0}=\{h_{1,0},\cdots,h_{n,0}\}$. Taking inspiration from the single-agent GRPO algorithm Guo et al. (2025); Liu et al. (2025b), at each turn $t$, each agent takes action by generating a group of responses $a^{\mathcal{G}}_{i,t}=\{a^{(1)}_{i,t},\cdots,a^{(G)}_{i,t}\}$ following its policy $\pi_{i}(\cdot|h_{i,t}^{\mathcal{G}})$ based on its observation-action history $h_{i,t}^{\mathcal{G}}=\{h^{(1)}_{i,t},\cdots,h^{(G)}_{i,t}\}$. The actions of individual agents are aggregated to form a group of joint actions $\mathbf{a}^{\mathcal{G}}_{t}=\{a^{\mathcal{G}}_{0,t},\cdots,a^{\mathcal{G}}_{n,t}\}$. The agents receive a group of joint rewards $r^{\mathcal{G}}_{t}$ for their responses $\mathbf{a}^{\mathcal{G}}_{t})$, which also conditions on the accessible part of the state $R(\cdot|s^{\text{acc},\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$. The joint actions triggers the transition $T(\cdot|s^{\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$, where agents receive new observations $o^{\mathcal{G}}_{i,t+1}=\{o^{(1)}_{i,t+1},\cdots,o^{(G)}_{i,t+1}\}$ and use them to construct histories $h^{\mathcal{G}}_{i,t+1}=\{h^{\mathcal{G}}_{i,t},a^{\mathcal{G}}_{i,t},o^{\mathcal{G}}_{i,t+1}\}$. This process continues until terminated at turn $H$.
@@
We employ stochastic gradient descent to train agents. Without explicit value models, estimating history-action values from a single rollout incurs high variance. To stabilize training, we estimate the expected return of the current state by averaging over a group of Monte Carlo samples $\{R^{(1)}_{t},\cdots,R^{(G)}_{t}\}$ for each branch $B$. As a result, we are able to generate a centralized estimate (which is common in MARL) without a large value model. For each turn $t$, the advantage of each joint action in the group is calculated as,
$\widehat{A}^{(g)}_{t}=R^{(g)}_{t}-\frac{1}{G}\sum_{g=1}^{|\mathcal{G}|}R^{(g)}_{t}$, (1)
@@
where $R^{(g)}_{t}=\sum_{\tau=t}^{H-1}r^{(g)}_{\tau}$. Inspired by GRPO Guo et al. (2025), Dr. GRPO Liu et al. (2025b), and MAPPO Yu et al. (2022), the centralized advantage values can be used to update policy $\pi_{i}$ (parameterized by $\theta_{i}$) for each agent $i$. MAGRPO does not have importance sampling and epsilon clipping since it is on-policy, and the weight of the KL divergence term is set to be 0 to encourage greater policy deviation from the base model,
@@
$J(\theta_{i})=\mathbb{E}_{\mathbf{o}_{0}\sim\mathcal{D},\mathbf{h}^{\mathcal{G}}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}}\Bigg[\frac{1}{|B|}\frac{1}{|\mathcal{G}|}\sum_{h_{i}^{\mathcal{G}}\in B}\sum_{g\in\mathcal{G}}\widehat{A}^{(g)}_{t}\log\pi_{\theta_{i}}(a_{i,t}^{(g)}|h_{i,t}^{\mathcal{G}})\Bigg],$ (2)
where the gradient is averaged across all the branches (states) and over the group of generations at $t$.
5 Experiments
(a) MAGRPO-ST on HE
(b) MAGRPO-MT on HE
(c) MAGRPO-ST on CHE
(d) MAGRPO-MT on CHE
Figure 2: Normalized returns: (a) Structural integrity (dashed grey); (b) Syntax correctness (dashed green); (c) Test pass rate (dashed red); (d) Cooperation (dashed yellow); (e) Total return (solid blue).
In large-scale software development, numerous developers cooperate to implement complex systems. Employing LLMs as developers is a promising direction, but coordinating them remains challenging due to diverse cooperation schemes and complex failures. In our experiments, we frame the coding collaboration task by having 2 Qwen2.5-Coder-3B agents to generate Python functions together, where a helper agent produces auxiliary functions to support a main function generator without direct communication. The outputs from both agents, along with the required libraries, are aggregated into complete code snippets.
5.1 Datasets
We first evaluate MAGRPO on the HumanEval (HE) dataset. HE comprises 164 handwritten programming problems, each containing a natural language description (prompt), a function signature (entry_point), and a set of unit tests (test). To guide learning, we design a level-based reward model that prioritizes fundamental aspects of code generation. Structural integrity verifies the presence and correctness of both main and auxiliary function definitions; syntactic correctness ensures compliance with Python syntax; test pass rate assesses functional correctness based on the proportion of successfully passed unit tests; and a cooperation quality bonus is granted when the main function properly invokes and utilizes the auxiliary function. The rewards are accumulated only when all requirements at each preceding level are satisfied.
However, most problems in HE are not designed for coding collaboration, as certain atomic operations (e.g., strlen(string)) can hardly be decomposed in a way that supports meaningful cooperation. These noisy instances bring instability into the training process or bias it toward inefficient cooperation schemes, such as the main agent merely wrapping the auxiliary function. Thus, we construct a cooperation-oriented code generation dataset, CoopHumanEval (CHE), for our evaluation. CHE includes both original HE problems with the potential for cooperation (e.g., prime_fib(n)) and additional handwritten programming problems (e.g., compare_areas(shapes)). Its data fields are the same as HE, with problem descriptions in prompt, function signatures in entry_point, and designed unit tests provided in test. The problems in CHE are readily decomposable, and agents can explore more effective cooperation schemes by training with this dataset. Both datasets are split into training and testing sets with a ratio of 25:8.
5.2 Baselines

### v3

@@ Abstract
A large amount of work has been done in Multi-Agent Systems (MAS) for modeling and solving problems with multiple interacting agents. However, most LLMs are pretrained independently and not specifically optimized for coordination. For example, existing LLM fine-tuning frameworks rely on individual rewards, which require complex reward designs for each agent to encourage collaboration. To address this challenge, we model LLM collaboration as a cooperative Multi-Agent Reinforcement Learning (MARL) problem. We develop a multi-agent, multi-turn algorithm, Multi-Agent Group Relative Policy Optimization (MAGRPO), to solve it, building on current RL approaches for LLMs as well as MARL techniques. Our experiments on LLM writing and coding collaboration demonstrate that fine-tuning multiple LLMs with MAGRPO enables agents to generate high-quality responses efficiently through effective cooperation. Our approach opens the door to using MARL methods for LLM collaboration and highlights the associated challenges.
Introduction
@@
State-of-the-art LLMs have demonstrated remarkable capabilities across diverse domains (llama3; gpt4; gemini). To adapt to specific applications or align with human preferences, fine-tuning has emerged as a critical training stage. Compared to supervised fine-tuning, Reinforcement Learning (RL) enables more generalizable learning for complex, multi-turn tasks through human-aligned reward design, making it an important technique for fine-tuning (ouyang; deepseek-r1; zieglerrlhf).
@@
Likewise, Multi-Agent Systems (MAS) have been extensively studied over the past decades, with substantial progress in modeling and solving problems involving multiple agents (weiss1999multiagent; van2008multi; yoavmas; peterstonemas). In particular, advances in cooperative MAS have demonstrated strong potential for enabling effective collaboration in distributed settings, such as games, robotics, and traffic control (smac; smacii; openai5; traffic; liutraffic; IJRR2025). These developments motivate the application of MAS principles and techniques to LLM collaboration, where multiple LLMs working together can solve more complex tasks more robustly and efficiently.
@@
There has been some recent work on coordinating multiple LLMs. Some approaches implement coordination at the inference stage, enabling agents to interact through debate, discussion, or verification (dudebate; autogen; duverifier). These methods operate at the prompt level, with fixed models that are not tuned toward coordination-centric objectives. The agents may have conflicting answers or spread incorrect information to other participants, limiting performance (whymultillmfail; debatefail). Moreover, the design of effective prompts remains difficult and unclear. Other approaches fine-tune agents independently with individual or role-conditioned rewards. However, they require carefully curated rewards for each individual or role (spiral; maft), and, as independent learning methods, lack convergence guarantees (Tan1993).
@@
We model LLM collaboration as a cooperative MARL problem (marl-book; marl) and formalize it as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (decpomdp). In LLM collaboration, multiple trainable LLMs generate responses synchronously based on their individual prompts. The external environment evolves according to the joint responses until the dialog ends. This general model allows a wide range of problems to be modeled and solved using versions of MARL algorithms. Following the efficient practice of Group Relative Policy Optimization (GRPO) (deepseek-math), we propose Multi-Agent GRPO (MAGRPO) that trains LLMs in a multi-turn setting. MAGRPO leverages centralized group-relative advantages for joint optimization, while preserving decentralized execution for each agent. The resulting method builds off of state-of-the-art LLM approaches in GRPO and MARL approaches for centralized training and decentralized execution, such as MAPPO (mappo). Our experiments demonstrate that MAGRPO is able to learn different cooperation schemes while producing efficient and high-quality responses.
@@
Our contributions can be summarized as follows: (i) We model the LLM collaboration as a cooperative MARL problem, where multiple LLMs cooperate to generate joint responses; (ii) We develop the MAGRPO algorithm, which optimizes agent cooperation through aligned rewards while maintaining decentralized execution to maintain efficiency; (iii) Our experiments demonstrate that fine-tuning with MAGRPO improves both response efficiency and quality in writing and coding collaboration; (iv) We provide an analysis of the limitations of existing approaches and outline open challenges in applying MARL to LLM collaboration.
Related Work
Test-Time Multi-Agent Interaction
@@
Recent work employs multiple agents with specialized roles interacting through diverse pipelines at test-time to enhance response quality. In multi-agent debate, agents iteratively formulate positions by reviewing other agents’ outputs, where the final decision or answer is determined by majority voting or a summarizer (dudebate; chandebate; liangdebate). Role-based approaches allocate tasks across specialized agents (autogen; chatdev; metagpt). An agent may function as a verifier to assess the correctness of outputs (skretaverifier; duverifier), or as a macro-planner to orchestrate workers’ responses. However, these multi-agent frameworks rely on prompt-level interactions among agents, often leading to ineffective communication and computational inefficiency. Moreover, the design of effective prompts and role assignment remains unclear, as prompts usually fail to reliably guide agent behavior, enforce role adherence, or support coherent coordination across tasks. These limitations motivate us to fine-tune LLMs in MAS to improve their cooperation.
Multi-Agent Fine-Tuning
@@
Recent work has explored fine-tuning LLMs to improve their performance across diverse domains, e.g., arithmetic reasoning, navigation, and hidden-role games (cory; amongus). These approaches typically employ individual rewards or rewards conditioned on specific roles (spiral; maft; marti2025; chen2025improvingretrievalaugmentedgenerationmultiagent). Such reward structures often require careful manual specification, and their underlying rationale is rarely well justified. The misaligned or conflicting incentives can hinder effective coordination. Moreover, these methods lack convergence guarantees, as each agent learns independently in a non-stationary environment where other agents are simultaneously updating their policies. In this paper, we focus on cooperative scenarios, where LLMs are trained with verified, human-aligned rewards.
Cooperative MARL for LLM Collaboration
@@
Since LLMs can act as a special class of agents, we leverage advances in MAS to improve their collaboration. We model LLM collaboration as a cooperative MARL problem and outline its unique challenges. We formalize this problem as a Dec-POMDP, as shown in Figure 1.
LLM Collaboration
LLM collaboration is the problem where LLMs cooperatively solve a class of tasks in MAS. Tasks are specified in language and provided to the LLMs as prompts. Each LLM generates a response synchronously in response to its own instructions. All responses jointly form a solution to the task.
@@
As discussed by a number of companies (nvidia_llm_agents; anthropic_multi_agent), a team of agents could generate a complex codebase. The code would be difficult, costly, and time-consuming to generate with a single agent, but a group of LLMs could do so quickly and cheaply. None of these agents is self-interested, but they can be trained using a scheme such as the one in this paper. Using a joint reward allows agents to specialize as needed to complete the task without complex prompt or reward engineering.
Problem Formalization
We formalize LLM collaboration as a subclass of the cooperative MARL problem, considering LLMs are agents and the types of problems they are solving. This problem is a form of Dec-POMDP (decpomdp), which allows cooperation through a joint reward while preserving scalable decentralized control. We show 2 instantiations of our framework in writing and coding tasks in the experiments.
@@
  • $\mathcal{S}$ denotes the full global state space. At turn $t$, a full state $s_{t}=(s_{t}^{\text{acc}},s_{t}^{\text{usr}})$ consists of parts that are accessible in the model and provided to the reward model $s_{t}^{\text{acc}}\in\mathcal{S}^{\text{acc}}$ (e.g., external models or systems), and the inaccessible user state $s_{t}^{\text{usr}}\in\mathcal{S}^{\text{usr}}$ that updates over time but is not maintainable. In a Dec-POMDP, the state can not be directly observed by the agents.
@@
In Dec-POMDP, since the states are not directly observed, each agent maintains its local observation-action history $\mathbf{h}=\{h_{1},\cdots,h_{n}\}$ to infer information about the state. A solution to a Dec-POMDP is a joint policy that maximizes the expected cumulative reward, $\boldsymbol{\pi}^{*}=\{\pi_{1}^{*},\cdots,\pi_{n}^{*}\}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{H-1}R(s^{\text{acc}}_{t},\mathbf{a}_{t})\right]$. A joint policy is a set of local policies $\pi_{i}$, which condition on the local observation-action history $h_{i,t}=\{o_{i,0},a_{i,0},\ldots,o_{i,t}\}$.
RL methods for Dec-POMDPs have become a popular topic (e.g., (COMA; MADDPG; Mordatch18; QMIX; QPLEX; mappo; marl-book; lyu2023centralized; Marchesini25)) with methods successful at scaling to large state, action, and observation spaces. Many methods use Centralized Training for Decentralized Execution (CTDE), where they use some centralized information during training (e.g., a centralized value function estimate) but are still able to execute in a decentralized manner when training is completed (CTDE).
Challenges in LLM Collaboration
@@
Representations in Natural Language
Unlike traditional cooperative MARL agents, LLM agents operate over natural language, receiving instructions and generating responses as sequences of tokens. MARL approaches could model this problem at the token or prompt/response level. At the token level, the number of actions and observations is smaller, but the problem horizon can be very long. At the prompt/response level, the actions and observations space is much larger, but the horizon is much shorter. Moreover, token-level rewards are often uninformative, as both queries and responses must form coherent and semantically meaningful structures. As adopted in prior RL methods (ouyang; deepseek-math; dpo), we model each LLM agent’s decision-making process as a direct mapping from input instructions to complete responses to enable efficient and stable training. Nevertheless, the best modeling and solution approaches remain an open question.
Training Paradigm
As mentioned above, many MARL methods use centralized training for decentralized execution (CTDE). Unfortunately, standard CTDE methods use centralized value models in the form of centralized critics (COMA; MADDPG; mappo) or mixers in value decomposition methods (QMIX; QPLEX), which allow additional information during training but do not scale well to very large action and observation spaces (such as those in LLM collaboration). Conversely, Decentralized Training and Execution (DTE) methods (marl) train a set of models, one for each agent in a decentralized manner, which are typically more scalable but do not use additional information during training (even when it is available). It is an open question which paradigm to use to maximize performance while maintaining scalability in the LLM collaboration problem. In this paper, we balance decentralized execution with centralized training using group-based Monte Carlo estimates. Experiments show the effectiveness of our approach on short-horizon tasks.
MAGRPO
@@
7:   Obtain joint rewards $r^{\mathcal{G}}_{t}$ from system
@@
12:   Estimate $\widehat{A}^{(g)}_{t}$, $\forall g\in\mathcal{G}$ according to Equation 1
13:   Calculate $J(\theta_{i})$, $\forall i\in\mathcal{I}$ according to Equation 2
@@
Algorithm 1 shows the procedure of MAGRPO. Given a dataset $\mathcal{D}$ containing task information (e.g., the descriptions of coding problems), $n$ LLMs that are optimized, each with a policy parameterized by $\theta_{i}$ and guided by a (shared) reward model $R$. In each episode, a task is sampled from the given dataset $\mathcal{D}$, which is used to construct initial observations $\mathbf{o}_{0}=\{o_{1,0},\cdots,o_{n,0}\}$ and histories $\mathbf{h}_{0}=\{h_{1,0},\cdots,h_{n,0}\}$. Taking inspiration from the single-agent GRPO algorithm (deepseek-math), at each turn $t$, each agent takes action by generating a group of responses $a^{\mathcal{G}}_{i,t}=\{a^{(1)}_{i,t},\cdots,a^{(G)}_{i,t}\}$ following its policy $\pi_{i}(\cdot|h_{i,t}^{\mathcal{G}})$ based on its observation-action history $h_{i,t}^{\mathcal{G}}=\{h^{(1)}_{i,t},\cdots,h^{(G)}_{i,t}\}$. The actions of individual agents are aggregated to form a group of joint actions $\mathbf{a}^{\mathcal{G}}_{t}=\{a^{\mathcal{G}}_{0,t},\cdots,a^{\mathcal{G}}_{n,t}\}$. The agents receive a group of joint rewards $r^{\mathcal{G}}_{t}$ for their responses $\mathbf{a}^{\mathcal{G}}_{t}$, which conditions on the accessible part of the state $R(\cdot|s^{\text{acc},\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$. The joint actions triggers the transition $T(\cdot|s^{\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$, where agents receive new observations $o^{\mathcal{G}}_{i,t+1}=\{o^{(1)}_{i,t+1},\cdots,o^{(G)}_{i,t+1}\}$ and use them to construct histories $h^{\mathcal{G}}_{i,t+1}=\{h^{\mathcal{G}}_{i,t},a^{\mathcal{G}}_{i,t},o^{\mathcal{G}}_{i,t+1}\}$. This process continues until terminated at turn $H$ or the task is completed.
@@
We employ stochastic gradient descent to train agents at the end of each episode. Without explicit value models, estimating history-action values incurs high variance. To stabilize training, we estimate the expected return of the current history by averaging over a group of Monte Carlo samples of the joint return $R^{\mathcal{G}}_{t}=\{R^{(1)}_{t},\cdots,R^{(G)}_{t}\}$. As a result, we are able to generate a centralized estimate (common in MARL) without a large value model. For each turn $t$, the advantage of each joint action in the group is calculated as,
$\widehat{A}^{(g)}_{t}=R^{(g)}_{t}-\frac{1}{G}\sum_{g=1}^{G}R^{(g)}_{t}$, (1)
@@
where $R^{(g)}_{t}=\sum_{\tau=t}^{H-1}r^{(g)}_{\tau}$. Inspired by GRPO (deepseek-math; drgrpo), and MAPPO (mappo), the centralized advantage values can be used to update policy $\pi_{i}$ (parameterized by $\theta_{i}$) for each agent $i$. MAGRPO does not have importance sampling and epsilon clipping since it is on-policy, and the KL divergence coefficient is set to be 0 to encourage greater policy deviation from the base model,
@@
$J(\theta_{i})=\mathbb{E}_{\mathbf{o}_{0}\sim\mathcal{D},\mathbf{h}^{\mathcal{G}}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}}\Bigg[\frac{1}{|G|}\sum_{g=1}^{G}\widehat{A}^{(g)}_{t}\log\pi_{\theta_{i}}(a_{i,t}^{(g)}|h_{i,t}^{\mathcal{G}})\Bigg].$ (2)
Experiments
We evaluate MAGRPO on LLM writing and coding collaboration. Datasets, reward specifications, and additional results are provided in the Appendix.
Writing Collaboration
Method Dataset Efficiency Article Quality (%) Return (%)
Speed Response Time Structure Consistency Coherence
Single Model TLDR 64.1 6.6 43.8 97.6 52.8 36.7
arXiv 65.4 6.5 51.2 87.2 71.1 44.9
Parallel Generation TLDR 185.6 2.1 25.9 98.3 56.5 23.2
arXiv 190.6 2.1 71.5 64.2 61.5 59.6
Sequential Generation TLDR 98.7 4.3 33.5 98.5 64.5 21.7
arXiv 85.8 4.3 92.4 97.8 64.3 87.7
One-Round Discussion TLDR 100.4 4.3 35.9 98.8 60.8 22.3
arXiv 95.4 4.3 84.6 71.8 66.0 76.6
MAGRPO (Ours) TLDR 202.3 2.1 98.7 97.1 78.5 94.5
arXiv 193.8 2.1 97.9 96.2 69.7 93.1
Table 1: Performance of MAGRPO against baselines on TLDR and arXiv. Speed (tokens/s) and response time (s) are measured on GeForce RTX 5090s. Results are normalized to the return scale. Bolds indicate the best performance on each dataset.
We explore LLM collaboration for article writing using MAGRPO across 2 classic tasks: summarization and expansion.
(a) TLDR summarization
(b) arXiv abstract expansion
Figure 2: Normalized returns on writing collaboration: (a) structural wellness (dashed green); (b) style consistency (dashed red); (c) coherence (dashed orange); (d) total rewards (solid blue). All returns are normalized to the return scale.
TLDR Summarization
When reading a long article, readers often seek to quickly grasp its core ideas. If the topic is of interest, they may wish to delve deeper into specific details while still avoiding a complete reading through the full document. This calls for a summarization system to generate summaries at varying levels of detail. We frame this task using TLDR summarization as an illustrative example.
The TLDR dataset comprises unabridged Reddit posts in the prompt and concise summaries appended by the author in the completion. In our experiment, 2 Qwen3-1.7B agents independently summarize the prompt without using completion. The first agent functions as a core-idea (TLDR) generator, producing a concise paragraph, while the second agent serves as a detailed summarizer, providing more comprehensive information.
To quantify the summarization quality, we employ a relatively simple combination of 3 metrics. Structure measures the lengths and the length ratio of the 2 summaries, to ensure the TLDR is concise and the detailed summary is sufficiently long. Style consistency is assessed using the normalized Jaccard similarity coefficient, calculated as the ratio of the intersection size to the union size of unique words (or n-grams) between responses. A high style consistency reward typically indicates that the summarizers adopt similar stylistic patterns while avoiding identical wording. Logical coherence is quantified by counting the occurrences of transition words. Positive reward is given for using transition words, but the reward decreases logarithmically as more are used. These metrics are simple approximations of what more complex reward models may evaluate. Other (simpler or more complex) metrics or reward models could also be used. The total reward combines these metrics through a weighted summation. More details regarding our reward model and hyperparameters are provided in the Appendix.
arXiv Expansion
Writing a long article typically requires contributions from multiple writers, each responsible for different sections. As a simple scenario, 2 agents can collaborate to generate introductions from the abstract of arXiv papers. The first agent outlines the research background and motivation, while the other presents the proposed methods and their experiments. The combined paragraphs should be coherent and consistent in style. Similar to the reward model in TLDR summarization, we employ the same evaluation metrics as proxies, with threshold hyperparameters specifically adjusted for this task.
Baselines